# Dielectric Tensor Prediction for Inorganic Materials Using Latent Information from Preferred Potential

### v1

@@ 5.3 Hyperparameters
The DTNN model contains 2 equivariant blocks (Fig. 3(a)). All the latent dimensions in MLP in Fig. 3(b) and (c) are set to 128 with a dropout rate of 0.2. The batch size is set to 64 during training. The AdamW optimizer[60] is utilized to optimize parameters with $\beta_{1}=0.9,\beta_{2}=0.999,\lambda=0.01$. The learning rate starts at $1\times 10^{-4}$ and a cosine annealing schedule is applied for tighter convergence. The gradient norm is clipped at 2.0 to stabilize the training. Early stopping is applied during training if no improvement on the validation accuracy is observed after 200 epochs.
@@ Acknowledgments
The authors thank So Takamoto for fruitful discussions and helpful comments. The authors thank Chikashi Shinagawa for assistance in DFPT computations using VASP.
@@
The materials with dielectric constants are freely accessible using The Materials Project API (https://next-gen.materialsproject.org/api).
@@
The code implementation of DTNN will be publicly available after submission to the journal.

### v2

@@ 5.3 Hyperparameters
The DTNet model contains 2 equivariant blocks (Fig. 3(a)). All the latent dimensions in MLP in Fig. 3(b) and (c) are set to 128 with a dropout rate of 0.2. The batch size is set to 64 during training. The AdamW optimizer [67] is utilized to optimize parameters with $\beta_{1}=0.9,\beta_{2}=0.999,\lambda=0.01$. The learning rate starts at $1\times 10^{-4}$ and a cosine annealing schedule is applied for tighter convergence. The gradient norm is clipped at 2.0 to stabilize the training. Early stopping is applied during training if no improvement on the validation accuracy is observed after 200 epochs.
5.4 Virtual Screening
We performed virtual screening based on iterative active learning to search for candidates for the two problems. The acquisition function uses an upper confidence bound (UCB) to balance exploration and exploitation during multi-round screening, defined by
$a(\chi;\kappa)=\mu(\chi)+\kappa\cdot\sigma(\chi)$ (28)
where $\chi$ is the input crystal structure, $\kappa$ is a hyperparameter controlling the trade-off and here $\kappa=1.5$ is selected for both tasks. $\mu(\chi)$ and $\sigma(\chi)$ represent the mean and standard deviation of polycrystalline dielectric constants and anisotropy ratios in the two respective tasks. In each task, we select the top 20 highest $a(x;\kappa=1.5)$ candidates with unit cell sizes of up to 20 atoms for validation by DFPT in each round in consideration of computational budget constraints. After each round, the dataset was updated with the newly validated results, and the next iteration was performed. In this case, finally, there were 60 new structures validated by DFPT computations across 3-round active learning for each discovery task.
5.5 DFPT Configuration
The density functional perturbation theory (DFPT) validation of the screened results was performed using the Vienna Ab initio Simulation Package (VASP) [68]. For the calculations, the generalized-gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) functional was employed, augmented by Hubbard U corrections with the same settings on the MP public documentation [69]. The projector augmented wave (PAW) method [61] was used to represent the core and valence electrons. The plane-wave basis set was truncated at an energy cutoff of 600 eV. The k-point density of 3000 k-points per reciprocal atom was used. A Gaussian smearing with a width of 0.05 eV was applied to account for partial occupancies in the electronic structure. A tight global electronic relaxation convergence criterion is applied with an energy difference tolerance of $1\times 10^{-7}$ eV. The electronic minimization utilizes a fairly robust mixture of the blocked-Davidson and RMM-DIIS algorithms to achieve efficient convergence. The precision setting was configured to "Accurate" to enhance the overall precision of the calculations.
@@ Acknowledgments
The authors thank So Takamoto for fruitful discussions and helpful comments. The authors thank Chikashi Shinagawa for assistance in DFPT computations using VASP. The research is supported by Preferred Networks Inc.
@@
The materials with dielectric constants are freely accessible using The Materials Project API (https://next-gen.materialsproject.org/api). Our preprocessed data from MP used in this study and our validated DFPT results are provided at https://github.com/pfnet-research/dielectric-pred.
@@
The implementation code of DTNet is available at https://github.com/pfnet-research/dielectric-pred. The pretrained model PFP is provided in the proprietary software named Matlantis. The code and trained parameters are not open-source. The related information of latest-version PFP can be found at https://matlantis.com/news/pfp-validation-for-public-v5-0-0.